what is the area of the grafh represented by |x|+|y|=1

Grade:12th pass

2 Answers

Shailendra Kumar Sharma
6 years ago
  1. st, Draw graph.
    For x>0, y>0, curve is x+y=1
    For x>0, y
    For xSo, area is (side)^2= 22.  On plotting these four straight lines, you'll see the final curve is a square of side
    For x
    0, curve is -x+y=1
Shailendra Kumar Sharma
6 years ago
First, draw graph.
For x>0, y>0, curve is x+y=1
For x
For x>0, y
For x> 
On plotting graph, you'll see its a square of side 2^1/2 ie root2. So, area = side^2= 2.
0, curve is -x+y=1